### v1

@@ Abstract
Non-cooperative interactions commonly occur in multi-agent scenarios such as car racing, where an ego vehicle can choose to overtake the rival, or stay behind it until a safe overtaking “corridor” opens. While an expert human can do well at making such time-sensitive decisions, the development of autonomous agents capable of rapidly reasoning about complex, potentially conflicting options is yet to be fully addressed. The recently developed nonlinear opinion dynamics (NOD) model shows promise in enabling fast (i.e., at an exponential rate) opinion formation and avoiding safety-critical deadlocks. However, it remains an open challenge to determine the model parameters of NOD automatically and adaptively, accounting for the ever-changing environment of interaction. In this work, we propose for the first time a learning-based, game-theoretic approach to synthesize a Neural NOD model from expert demonstrations, given as a dataset containing (possibly incomplete) state and action trajectories of interacting agents. The learned NOD can be used by existing dynamic game solvers to plan decisively while accounting for the predicted change of other agents’ intents, thus enabling situational awareness in planning. We demonstrate Neural NOD’s ability to make fast and robust decisions in a simulated autonomous racing example, leading to tangible improvements in safety and overtaking performance over state-of-the-art data-driven game-theoretic planning methods.
Keywords:
nonlinear opinion dynamics inverse games neural synthesis
1 Introduction
As autonomous multi-agent systems evolve toward an unprecedented complexity, especially in high-stakes scenarios such as car racing, the necessity for reliable decision-making in real-time becomes paramount. Rapid decision-making is critical in these settings, not only for competitiveness but also for ensuring safety during complex, close-proximity interactions. For example, in autonomous car racing, the ego vehicle must decide, in a split second, between overtaking a rival car without crashing into the rival or strategically trailing it to capitalize on emerging opportunities. While existing game-theoretic approaches [5] effectively capture non-cooperative interactions, they often fall short in real-time calibration of strategies to account for the evolving interactions with other agents. Although multi-agent reinforcement learning (RL) and imitation learning approaches, such as [47, 53], show promise in performance and scalability, they may require massive amounts of data to mitigate generalization issues [43], and it remains unclear how these methods can ensure fast adaptation to changing environments and avoid safety-critical indecisions due to lack of interpretability in decision-making.
@@ 1 Introduction
The recently developed NOD model [6] provides a principled framework to describe opinion exchanges among agents and guarantees indecision breaking. More specifically, an opinion state numerically represents agents’ agreement or disagreement on a topic that is relevant to collective decision-making, thus being a natural mathematical tool for modeling multi-robot non-cooperative interactions. NOD achieves rapid opinion formation with nonlinear bifurcation, at which point the process is ultrasensitive to external inputs, e.g., physical interactions among agents. However, it remains an open challenge to select suitable NOD model parameters such that the ego agent can reason how ongoing interactions may drive rapid changes in its own situational awareness and intended course of action as well as those of other agents.
Contributions and Paper Organization. We propose for the first time an inverse-game-based approach to synthesize an NOD model parameterized by a deep neural network (DNN). This model is used for automatic cost tuning by online dynamic game solvers to enable time-sensitive multi-agent decision-making. We provide verifiable conditions under which the model is guaranteed to avoid indecisions. We deploy the Neural NOD learned from real human racing data for simulated autonomous car racing and compare our approach against the state-of-the-art data-driven game-theoretic planning baseline.
The paper is structured as follows. Section 2 discusses related work in multi-agent decision-making. Section 3 reviews dynamic games and NOD. In Section 4, we discuss how to learn a Neural NOD model from data leveraging inverse dynamic games, and use it for split-second decision-making in multi-agent settings. Section 5 presents simulation results in autonomous car racing. We discuss limitations and future work in Section 6. Finally, Section 7 concludes the paper.
2 Related Work
@@
Multi-Agent Motion Planning. Early efforts in multi-agent motion planning focus on developing decentralized or distributed algorithms, leveraging single-agent planning approaches such as graph search [46], model predictive control [12], and trajectory optimization [1]. However, these methods can struggle with general non-cooperative interactions involving self-interested, potentially competing agents. Game-theoretic motion planning explicitly reasons about the coupled interests among agents, and has been widely used for multi-agent non-cooperative interactions, e.g., autonomous driving [44, 13, 50, 24], human–robot interaction [31, 4, 21], and distributed control systems [35, 20, 51]. Computing equilibria is oftentimes computationally intractable for general dynamic games due to their inherent non-convexity involving, e.g., nonlinear system dynamics, non-convex cost functions, and collision-avoidance constraints. To tackle such computation challenges, Fridovich-Keil et al. [14] apply iterative linear-quadratic (ILQ) approximation to solve general-sum dynamic games, enabling real-time computation of approximate Nash equilibria. Recent work [32] extends this approach to compute a mixed Nash strategy that blends task optimality with a multi-modal data-driven reference policy.
@@
In contrast to model-based optimization, data-driven methods have also shown competitive performance in interactive autonomy. GT Sophy [52] demonstrates that a well-trained deep RL policy can win a head-to-head competition against some of the world’s best drivers in Gran Turismo, a popular car racing video game. However, the policy does not explicitly reason the strategic interactions among players. To investigate this gap, Chen et al. [10] combine RL with self-supervised learning that models the opponent behaviors, and show that the resulting policy outperforms methods with manually designed opponent strategies in multi-agent car racing. Song and Lin et al. [47] use curriculum learning to improve the overtaking performance in Gran Turismo races. Recent work [23] leverages belief-space planning to actively reduce opponent uncertainty in adversarial interactions, and uses deep RL to synthesize a best-effort reach-avoid policy. Despite promising performance and scalability, these black-box methods lack game-theoretic interpretations when deployed for complex interactions, and may suffer from generalization issues when data are insufficient [43].
Our proposed motion planning framework combines the best of both worlds by offering a game-theoretically meaningful solution with a model-based solver, which is guided by a Neural NOD model learned from interaction data for fast and decisive motion planning in rivalrous, safety-critical interactions.
Inverse Dynamic Games. Unlike forward games that focus on computing equilibria strategies, inverse dynamic games are useful when information of the game, e.g., cost parameters, is unknown or unspecified, but may be learned from data. Peters et al. [42] leverages differentiable optimization [3] to learn players’ a priori unknown objectives of a non-cooperative dynamic game based on noise-corrupted, partial state observations. The optimization couples past state estimates with future state predictions, thus enabling simultaneous online learning, prediction, and decision-making. Liu et al. [33] further extend this idea to solve constrained inverse dynamic games, which yield generalized equilibria strategies. Another concurrent follow-up work [29] proposes an inverse dynamic game framework assuming all players adopt a feedback Nash equilibrium strategy. However, these inverse game methods mainly focus on learning static cost parameters from a single trajectory. As such, the inverse game needs to be re-solved when the environment (e.g., the initial state) has evolved, posing computational challenges at runtime.
Our approach builds on existing inverse game approaches to learn a Neural NOD model, and use it in forward game solvers to facilitate time-sensitive decision-making in rapidly evolving interactive environments. To the best of our knowledge, our work is the first to apply inverse dynamic games to learn a deep model from many trajectories, which practically generalizes to arbitrary initial states at deployment time.
@@
Nonlinear Opinion Dynamics. NOD is pioneered by Bizyaeva et al. [6] for deadlock-free decision-making in multi-agent interactions. Cathcart et al. [9] use NOD to break social deadlocks in a human–robot corridor passing problem. Amorim et al. [2] propose an NOD-based decision-making framework for robots choosing between different spatial tasks while adapting to environmental conditions, where opinions represent agents’ preference for a particular task. Paine et al. [39] use NOD to make group decisions for autonomous multi-robot systems, demonstrating robust and interpretable collective behaviors in a field test involving multiple unmanned surface vessels. Hu et al. [22] design for the first time a general algorithm to automatically synthesize an NOD model. They propose to construct NOD parameters based on dynamic game value functions, thus making the opinion evolution dependent on agents’ physical states. However, this approach assumes that agents are facing a set of mutually exclusive options, and it cannot effectively integrate data-driven prior knowledge to facilitate online decision-making. For a comprehensive review of NOD, we refer the readers to Leonard et al. [28].
@@
The proposed Neural NOD can be trained over a dataset of diverse interaction trajectories, account for general opinion spaces, and is plug-and-play for any model-based game solver. Compared to previous NOD with static model parameters, the Neural NOD dynamically adjusts its parameters according to the evolving physical states, allowing the ego robot to rapidly adapt its decision to the fast-changing environment.
3 Preliminaries
@@
where $c^{i}_{k}(\cdot)$ is the stage cost, $\theta^{i}\in\mathbb{R}^{n_{\theta^{i}}}$ is the (possibly unknown) parameter of the stage cost, and ${\pi}:=({\pi}^{1}(\cdot),\ldots,{\pi}^{N_{a}}(\cdot))$ is a tuple containing all players’ control policies, which determines the information structure [5, Ch. 5] of the game, e.g., open-loop if ${u}^{i}_{t}={\pi}^{i}({x}_{0})\in\mathbb{R}^{m_{i}}$ and feedback if ${u}^{i}_{t}={\pi}^{i}({x}_{t})\in\mathbb{R}^{m_{i}}$. Policy tuple ${\pi}$ can constitute different equilibrium types of the game, among which the most common ones are Nash [37] and Stackelberg [48]. In this paper, we do not adhere to a specific information structure or equilibrium type.
Inverse Dynamic Games. When cost parameters $\theta:=(\theta^{1},\ldots,\theta^{N_{a}})$ are initially unknown, we can solve an inverse dynamic game [42, 29, 33] to identify them from trajectory data. This is formulated as a maximum likelihood estimation (MLE) problem:
@@
where $p(\mathbf{{y}}\mid{\mathbf{x}},{\mathbf{u}})$ is the likelihood of observed trajectory data $\mathbf{{y}}:={y}_{[0:T]}$ given state trajectories ${\mathbf{x}}:={x}_{[0:T]}$ and control sequences ${\mathbf{u}}:={u}_{[0:T-1]}$ that correspond to a (user-specified) equilibrium, and $\Gamma(\theta)$ is the set of all such equilibrium solutions parameterized by $\theta$. MLE (3) can handle corrupted data: the initial state ${x}_{0}$ is not assumed to be known, and the observation data $\mathbf{{y}}$ may be noisy and missing certain entries (e.g., at specific time indices). An inverse game routine such as [42, 29, 33] solves MLE (3) by applying gradient ascent to update the value of $\theta$, where gradient ${\nabla}_{\theta}p(\cdot)$ is computed by differentiating through a forward game solver that produces an equilibrium solution $({\mathbf{x}},{\mathbf{u}})\in\Gamma(\theta)$. Existing inverse game methods predominantly learn a static, context-agnostic cost parameter $\theta$ along a single game trajectory ${\mathbf{x}}$. In Section 4, we extend established inverse game techniques to learn an adaptive, context-aware cost parameter governed by an NOD from a dataset containing diverse gameplay trajectories.
Nonlinear Opinion Dynamics. The NOD model [6] enables fast and flexible multi-agent decision-making. Consider multi-agent system (1), in which each agent ${i}$ is faced with an arbitrarily large (but finite) number of ${N_{o^{i}}}$ options. For every ${i}\in\mathcal{I}_{a}$ and $\ell\in\mathcal{I}_{o^{i}}:=\{1,\ldots,{N_{o^{i}}}\}$, define ${z}^{{i}}_{\ell}\in\mathbb{R}$ to be the opinion state of agent ${i}$ about option $\ell$. The more positive (negative) is ${z}^{{i}}_{\ell}$, the more agent ${i}$ favors (disfavors) option $\ell$. We say agent ${i}$ is neutral (i.e., undecided) about option $\ell$ if ${z}^{{i}}_{\ell}=0$. Compactly, we define ${z}^{i}:=({z}^{i}_{1},\ldots,{z}^{i}_{{N_{o^{i}}}})$, and ${z}:=({z}^{1},\ldots,{z}^{{N_{a}}})$ as the opinion state of agent $i$ and the joint system, respectively. The evolution of opinion state in continuous time is governed by the NOD model, which is an ordinary differential equation:
@@
${S}^{{i}}_{\ell}(\cdot)={S}_{1}\Big{(}\alpha^{{i}}_{\ell}{z}^{{i}}_{{\ell}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)}$
which satisfies for $r\in\{1,2\}$, ${S}_{r}(0)=0$, ${S}^{\prime}_{r}(0)=1$, ${S}^{\prime\prime}_{r}(0)\neq 0$, and ${S}^{\prime\prime\prime}_{r}(0)\neq 0$. Valid choices for ${S}_{r}(\cdot)$ include the sigmoid function and the hyperbolic tangent function $\tanh$. In NOD model (4), $D^{{i}}=\operatorname{diag}({d}^{i}_{1},\ldots,{d}^{i}_{N_{o^{i}}})$ is the damping matrix with each ${d}^{i}_{\ell}>0$, ${b}^{i}$ represents agent ${i}$’s own bias, ${\lambda}>0$ is the attention value on nonlinear opinion exchange, which is shared across all agents, $\alpha^{{i}}_{\ell}\geq 0$ is the self-reinforcement gain, $\beta^{{i}}_{\ell{p}}\geq 0$ is the same-agent inter-option coupling gain, $\gamma^{{i}{j}}_{\ell}$ is the gain of the same-option inter-agent coupling with another agent ${j}$, and $\delta^{{i}{j}}_{\ell{p}}$ is the gain of the inter-option inter-agent coupling with another agent ${j}$. In order to guide a dynamic game with NOD, we consider the discrete-time variant of the NOD model jointly for all ${i}\in\mathcal{I}_{a}$:
@@
which may be obtained by applying time discretization (e.g., forward Euler or Runge-Kutta method) to the continuous-time NOD model (4). While NOD has demonstrated efficacy in multi-agent decision-making, automatic synthesis of its parameters that can change adaptively based on physical states, e.g., $\alpha^{i}_{t}({x}_{t})$, largely remains an open problem. In the next section, we propose a novel inverse-game-based approach that learns a general DNN-parameterized NOD model, which allows the opinion state ${z}$ to be influenced by physical states ${x}$.
Remark 1 (NOD Tunable Sensitivity)
The NOD provides tunable speed and sensitivity, that is, a “fast” decision is governed by a small number of tunable parameters. Therefore, NOD yields rapid decision-making only if it is needed by the robot’s specific task. This trait is inherent in the NOD model. Alternative approaches, such as Bayesian methods, may require additional (manual) modifications to the model on a case-by-case basis to alter its sensitivity to external inputs. In this paper, we will learn a DNN for automatic, context-aware tuning of NOD parameters.
4 Neural NOD for Interactive Robotics

### v2

@@ Abstract
Non-cooperative interactions commonly occur in multi-agent scenarios such as car racing, where an ego vehicle can choose to overtake the rival, or stay behind it until a safe overtaking “corridor” opens. While an expert human can do well at making such time-sensitive decisions, autonomous agents are incapable of rapidly reasoning about complex, potentially conflicting options, leading to suboptimal behaviors such as deadlocks. Recently, the nonlinear opinion dynamics (NOD) model has proven to exhibit fast opinion formation and avoidance of decision deadlocks. However, NOD modeling parameters are oftentimes assumed fixed, limiting their applicability in complex and dynamic environments. It remains an open challenge to determine such parameters automatically and adaptively, accounting for the ever-changing environment. In this work, we propose for the first time a learning-based and game-theoretic approach to synthesize a Neural NOD model from expert demonstrations, given as a dataset containing (possibly incomplete) state and action trajectories of interacting agents. We demonstrate Neural NOD’s ability to make fast and deadlock-free decisions in a simulated autonomous racing example. We find that Neural NOD consistently outperforms the state-of-the-art data-driven inverse game baseline in terms of safety and overtaking performance.
@@
I Introduction
As autonomous multi-agent systems evolve toward an unprecedented complexity, especially in high-stakes scenarios such as car racing, the necessity for reliable decision-making in real-time becomes paramount. Rapid decision-making is critical in these settings, not only for performance but also for ensuring safety during complex, close-proximity interactions. For example, in autonomous car racing (Figure 1), the ego vehicle must decide, in a split second, between overtaking a rival car without crashing into the rival or strategically trailing it to capitalize on emerging opportunities.
@@ I Introduction
Game-theoretic motion planning techniques explicitly reason about the coupled interests among agents, and have been widely used for multi-agent non-cooperative interactions, e.g., autonomous driving [schwarting2019social, wang2021game], human–robot interaction [li2019differential, bajcsy2021analyzing, hu2023activeIJRR], and distributed control systems [maestre2011distributed, hu2020non, williams2023distributed]. Computing equilibria for general dynamic games is oftentimes computationally intractable due to the inherent non-convexity involving, e.g., nonlinear system dynamics, non-convex cost functions, and collision-avoidance constraints. To tackle such computation challenges, Fridovich-Keil et al. [fridovich2020efficient] apply iterative linear-quadratic (ILQ) approximations to solve general-sum dynamic games, enabling real-time computation of approximate Nash equilibria. Recent work [lidard2024blending] extends this approach to the stochastic setting by blending a game policy (with a fixed payoff) with a data-driven reference policy. While these game-theoretic approaches can effectively capture non-cooperative interactions, they often fall short in rapid calibration of strategies to account for agents’ evolving intents.
@@
As an alternative to model-based optimization, data-driven methods have also shown competitive performance in interactive autonomy. GT Sophy [wurman2022outracing] demonstrates that a well-trained deep reinforcement learning (RL) policy can win a head-to-head competition against some of the world’s best drivers in Gran Turismo, a popular car racing video game. However, the policy does not explicitly reason about the strategic interactions among players. To investigate this gap, Chen et al. [chen2023learn] combine RL with self-supervised learning that models the opponent behaviors and show that the resulting policy outperforms methods with manually designed opponent strategies in multi-agent car racing. Recent work [hu2023belgame] leverages belief-space safety analysis to actively reduce opponent uncertainty in adversarial interactions. Despite promising performance and scalability, these black-box methods may lack strategic reasoning when deployed for complex interactions and can struggle with practical limitations such as poor generalization, particularly when faced with insufficient or unrepresentative data [ren2021generalization].
The recently developed NOD model [bizyaeva2022nonlinear] provides a principled way to make deadlock-free decisions in multi-agent interactions. Opinion states numerically represent agents’ agreement or disagreement on a decision. NOD achieves rapid opinion formation through a nonlinear bifurcation. A local bifurcation is a change in the number and/or stability of equilibrium solutions as a parameter varies through a bifurcation point, which corresponds to a singularity in the dynamics. Near this point, the process is ultrasensitive to external inputs, e.g., physical interactions among agents, and the implicit threshold on the input associated with rapid opinion formation can be tuned by modulating parameters [leonard2024fast].
@@
Cathcart et al. [cathcart2022opinion] use NOD to break social deadlocks in a human–robot navigation problem. Amorim et al. [amorim2024threshold] propose an NOD-based decision-making framework for robots choosing between different spatial tasks while adapting to environmental conditions, where opinions represent agents’ preference for a particular task. Paine et al. [Paine2024] leverage NOD to make group decisions for autonomous multi-robot systems, demonstrating robust and interpretable collective behaviors in a field test involving multiple unmanned surface vessels. Hu et al. [hu2023emergent] design for the first time a general algorithm to automatically synthesize an NOD model. They propose to construct NOD parameters based on dynamic game value functions, thus making the opinion evolution dependent on agents’ physical states. However, this approach assumes that agents face a set of mutually exclusive options, and it cannot effectively integrate data-driven prior knowledge to facilitate online decision-making. For a comprehensive review of NOD, we refer the readers to Leonard et al. [leonard2024fast]. For general multi-agent non-cooperative interaction, it remains an open challenge to select suitable NOD model parameters such that the ego agent can reason how ongoing interactions may drive rapid changes in its intended course of action as well as those of other agents.
@@
Contributions. We propose for the first time a learning-based and game-theoretic approach to synthesize a Neural NOD model for time-sensitive multi-agent decision-making. The model can be trained over a dataset of diverse interaction trajectories, can account for general opinion spaces, and can be plug-and-play used by any model-based game solver for automatic cost tuning. Compared to previous NOD with static model parameters, the Neural NOD dynamically adjusts its parameters according to the evolving physical states, allowing the ego robot to rapidly adapt its decision to the fast-changing environment. We provide verifiable conditions under which the model is guaranteed to avoid indecision. We deploy the Neural NOD learned from real human racing data for simulated car racing and compare our approach with the state-of-the-art data-driven game-theoretic planning baseline.
II Preliminaries
@@
where $c^{i}_{k}(\cdot)$ is the cost of stage $k$, $\theta^{i}\in\mathbb{R}^{n_{\theta^{i}}}$ is the (possibly unknown) parameter of the stage cost, and ${\pi}:=({\pi}^{1}(\cdot),\ldots,{\pi}^{N_{a}}(\cdot))$ is a tuple containing all players’ control policies, which determines the information structure [bacsar1998dynamic, Ch. 5] of the game, e.g., open-loop if ${u}^{i}_{t}={\pi}^{i}({x}_{0})\in\mathbb{R}^{m_{i}}$ and feedback if ${u}^{i}_{t}={\pi}^{i}({x}_{t})\in\mathbb{R}^{m_{i}}$. Policy tuple ${\pi}$ can constitute different equilibrium types of the game, among which the most common ones are Nash [nash1951non] and Stackelberg [stackelberg1934]. In this paper, we do not adhere to a specific information structure or equilibrium type.
Inverse Dynamic Games. When cost parameters $\theta:=(\theta^{1},\ldots,\theta^{N_{a}})$ are initially unknown, we can solve an inverse dynamic game [peters2023online, li2023cost, liu2023learning] to identify these parameters from trajectory data. This is formulated as a maximum likelihood estimation (MLE) problem:
@@
where $p(\mathbf{{y}}\mid{\mathbf{x}},{\mathbf{u}})$ is the likelihood of observed trajectory data $\mathbf{{y}}:={y}_{[0:T]}$ given state trajectories ${\mathbf{x}}:={x}_{[0:T]}$ and control sequences ${\mathbf{u}}:={u}_{[0:T-1]}$ that correspond to a (user-specified) equilibrium, and $\Gamma(\theta)$ is the set of all such equilibrium solutions parameterized by $\theta$. MLE (3) can handle corrupted data: the initial state ${x}_{0}$ is not assumed to be known, and the observation data $\mathbf{{y}}$ may be noisy and missing certain entries (e.g., at specific time indices). An inverse game routine such as [peters2023online, li2023cost, liu2023learning] solves MLE (3) by applying gradient ascent to update the value of $\theta$, where gradient ${\nabla}_{\theta}p(\cdot)$ is computed by differentiating through a forward game solver that produces an equilibrium solution $({\mathbf{x}},{\mathbf{u}})\in\Gamma(\theta)$.
Nonlinear Opinion Dynamics. The NOD model [bizyaeva2022nonlinear] enables fast and flexible multi-agent decision-making. Consider multi-agent system (1), in which each agent ${i}$ is faced with an arbitrarily large (but finite) number of ${N_{o^{i}}}$ options. For every ${i}\in\mathcal{I}_{a}$ and $\ell\in\mathcal{I}_{o^{i}}:=\{1,\ldots,{N_{o^{i}}}\}$, define ${z}^{{i}}_{\ell}\in\mathbb{R}$ to be the opinion state of agent ${i}$ about option $\ell$. The more positive (negative) is ${z}^{{i}}_{\ell}$, the more agent ${i}$ favors (disfavors) option $\ell$. We say agent ${i}$ is neutral (i.e., undecided) about option $\ell$ if ${z}^{{i}}_{\ell}=0$. Compactly, we define ${z}^{i}:=({z}^{i}_{1},\ldots,{z}^{i}_{{N_{o^{i}}}})$, and ${z}:=({z}^{1},\ldots,{z}^{{N_{a}}})$ as the opinion state of agent $i$ and the joint system, respectively. The evolution of opinion state in continuous time is governed by the NOD model adapted from [bizyaeva2022nonlinear]:
@@
${S}^{{i}}_{\ell}(\cdot)=$ ${S}_{1}\Big{(}\alpha^{{i}}_{\ell}{z}^{{i}}_{{\ell}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\gamma^{{i}{j}}_{\ell}{z}^{{j}}_{{\ell}}\Big{)}$
$+\sum_{{p}\in\mathcal{I}_{\theta_{i}}\setminus\{{\ell}\}}{S}_{2}\Big{(}\beta_{\ell{p}}^{{i}}{z}^{{i}}_{{p}}+\textstyle\sum_{{j}\in\mathcal{I}_{a}\setminus\{{i}\}}\delta^{{i}{j}}_{\ell{p}}{z}^{{j}}_{{p}}\Big{)},$
which satisfies for $r\in\{1,2\}$, ${S}_{r}(0)=0$, ${S}^{\prime}_{r}(0)=1$, ${S}^{\prime\prime}_{r}(0)\neq 0$, and ${S}^{\prime\prime\prime}_{r}(0)\neq 0$. Valid choices for ${S}_{r}(\cdot)$ include the sigmoid function and the hyperbolic tangent function $\tanh$. In (4), $D^{{i}}=\operatorname{diag}({d}^{i}_{1},\ldots,{d}^{i}_{N_{o^{i}}})$ is the damping matrix with each ${d}^{i}_{\ell}>0$, ${b}^{i}$ represents agent ${i}$’s own bias, ${\lambda}>0$ is the attention value on nonlinear opinion exchange, which here taken to be shared across all agents, $\alpha^{{i}}_{\ell}\geq 0$ is the self-reinforcement gain, $\beta^{{i}}_{\ell{p}}\geq 0$ is the same-agent inter-option coupling gain, $\gamma^{{i}{j}}_{\ell}$ is the gain of the same-option inter-agent coupling with another agent ${j}$, and $\delta^{{i}{j}}_{\ell{p}}$ is the gain of the inter-option inter-agent coupling with another agent ${j}$. In order to guide a dynamic game with NOD, we consider the discrete-time version of the NOD model jointly for all ${i}\in\mathcal{I}_{a}$:
@@
which may be obtained by applying time discretization (e.g., forward Euler or Runge-Kutta method) to the continuous-time NOD model (4). While NOD has demonstrated efficacy in multi-agent decision-making, automatic synthesis of its parameters that can change adaptively based on physical states, e.g., $\alpha^{i}_{t}({x}_{t})$, largely remains an open problem. In the next section, we propose a novel inverse-game-based approach that learns a general deep neural network (DNN)-parameterized NOD model, which allows the opinion state ${z}$ to be influenced by physical states ${x}$.
III Neural NOD for Interactive Robotics